f(X)=x+8; g(x)=7x-3
find (f+g)(X)

Found 2 solutions by stanbon, It is costly:
Answer by stanbon(75887)

f(X)=x+8; g(x)=7x-3
find (f+g)(X) = f(x) + g(x) = (x+8)+(7x-3) = 8x+5
------------------
Cheers,
Stan H.
------------

Answer by It is costly(175)

(f+g)(x)=x+8+7x-3
(f+g)(x)=8x+5